# Adaptive Local Combining with Decentralized Decoding for Distributed Massive MIMO

### v2

@@ IV-B Performance Evaluation
Fig. 3 shows the distribution of strong pilot decisions ($L_{\mathcal{S}_{m}}$) across all APs. The baseline PFZF and PWPFZF schemes exhibit a pronounced peak at $L_{\mathcal{S}_{m}}{=}6$-77, revealing their rigid threshold-based strategy that sacrifices excessive degrees of freedom for interference suppression regardless of local conditions. In striking contrast, the proposed G-PFZF and G-PWPFZF schemes show fundamentally different behavior: G-PFZF peaks at 2-4 strong pilots while G-PWPFZF shows an even more conservative distribution peaking at 1-3. This efficient allocation preserves spatial resources for signal enhancement. Most notably, the significant non-zero count at $L_{\mathcal{S}_{m}}{=}0$ demonstrates our algorithm’s ability to adaptively default to simple MR combining when it is the optimal strategy for a given AP’s local conditions. This demonstrates a seamless, adaptive switching between local zero-forcing and local MR processing that is impossible for the threshold-bound baselines.
@@
Fig. 4 depicts the sum SE versus the number of UEs ($T$) for various combining schemes. The proposed decentralized optimization framework delivers significant and scalable performance gains, with G-PFZF outperforming its baseline PFZF by 6.5% to 9.5% and G-PWPFZF surpassing PWPFZF by 5.5% to 10%, with the gaps widening as the number of UEs increases. Better performance of the proposed schemes highlights the critical limitation of the baseline PFZF and PWPFZF schemes, which rely on a hard, network-wide threshold using LSFCs to rigidly classify users as strong or weak at every AP. In contrast, the proposed framework introduces a decentralized optimization algorithm, executed independently by each AP, to determine a better pilot-based grouping strategy. This allows an AP to adaptively select none, a subset, or all of the pilots as strong, based on its local channel conditions and specific geometric configuration. This per-AP flexibility enables a more efficient and dynamic trade-off between utilizing degrees of freedom for interference suppression and desired signal enhancement. Consequently, the proposed G-PWPFZF scheme achieves better performance over the G-PFZF scheme by suppressing interference of the weak UEs from the strong UEs at the APs.
@@
Fig. 5 illustrates the sum SE versus the number of orthogonal pilots $L_{p}$, demonstrating the improved performance of the proposed G-PFZF and G-PWPFZF schemes over their respective baseline counterparts. The proposed schemes achieve significant performance improvements of 6-14% for G-PFZF and 6-13% for G-PWPFZF compared to their respective baselines across different pilot lengths, with the performance gap widening as the number of pilots increases. This expanding gap highlights the advantage of the proposed adaptive user grouping strategy over the rigid network-wide threshold approach used in baseline schemes. The baseline schemes, must rigidly limit the number of strong pilots to avoid violating the fundamental condition for zero-forcing ($L_{\mathcal{S}_{m}}{>}A$), which leads to inefficient interference management. In contrast, our proposed decentralized framework empowers each AP to independently and optimally select its set of strong pilots, inherently ensuring that the selection is both locally optimal and always feasible ($L_{\mathcal{S}_{m}}{<}A$). This inherent adaptability allows our methods to leverage larger pilot sets more effectively, maximizing array gain and suppressing interference, which explains the observed widening of the performance gap with the number of pilots.
Figure 6: The uplink sum SE comparison for various combining schemes.
Fig. 6 illustrates the sum SE versus the number of antennas per AP ($A$) for various combining schemes, revealing a characteristic performance pattern that highlights the adaptability of our proposed framework. The G-PFZF and G-PWPFZF schemes demonstrate consistent superiority across all antenna configurations, with the performance gap exhibiting a distinctive trajectory that reflects adaptive resource utilization. At the minimal antenna configuration (2 antennas), where spatial degrees of freedom $A-L_{\mathcal{S}_{m}}$ are severely constrained, both proposed and baseline schemes face fundamental challenges in interference suppression. Nevertheless, our adaptive grouping strategy achieves measurable gains of 2.5% for G-PFZF and 6% for G-PWPFZF through an adaptive selection of combining schemes, dynamically switching between the G-PFZF/G-PWPFZF, and the MR at each AP based on local conditions. In the mid-range antenna configuration (4-8 antennas), where careful trade-off between interference suppression and signal enhancement becomes crucial, our proposed method demonstrates performance gain of 10-14% for G-PFZF and 10-18% for G-PWPFZF, showcasing their superior grouping capabilities. As antenna numbers increase further (10-16 antennas), providing abundant spatial resources that benefit all schemes, the performance gap narrows to 4.5-6.5% for G-PFZF and 3.5-5.5% for G-PWPFZF, though our approach maintains consistent superiority.
Figure 7: The uplink sum SE comparison for various combining schemes.
Fig. 7 illustrates the sum SE versus the number of UEs ($T$), comparing local and LSFD-based combining schemes. For our G-PFZF and G-PWPFZF schemes, the performance gap between local weights and optimal LSFD remains minimal at just 3% across all UE densities. This marginal performance difference is dramatically outweighed by the substantial system benefits: our local weight approach eliminates the need for complex network-wide coordination, reducing computational cost at the CPU by avoiding large-scale matrix inversions and reducing fronthaul overhead by eliminating the need to share channel state information across the network. Despite this streamlined architecture, our locally-weighted G-PFZF and G-PWPFZF schemes still outperform the baseline PFZF and PWPFZF schemes with LSFD by margins of 4.5-9% and 3.5-7.5% respectively. This demonstrates that the gains from our adaptive per-AP grouping strategy are so substantial that they outperform the benefits of optimal coordination applied to suboptimal threshold-based grouping. These results confirm that our fully decentralized approach achieves a better balance between performance and practicality.
Figure 8: The uplink sum SE comparison for various combining schemes.
Fig. 8 illustrates the sum SE versus the number of pilots ($L_{p}$), comparing local and LSFD-based combining schemes. The performance gap between local and LSFD-based weights is less than 5.5% for sparse pilot configurations, where high pilot contamination makes centralized LSFD better equipped to handle interference. However, as the number of pilots increases to moderate and high values, the gap reduces to less than 1%, demonstrating near-identical performance. This trend occurs because with sufficient orthogonal pilots, both approaches have similar capability to mitigate pilot contamination. Notably, even in the worst-case scenario with limited pilots, the performance penalty remains modest, validating that the substantial fronthaul and complexity reductions of our decentralized decoding approach come with acceptable performance trade-offs across all pilot sequences. Furthermore, the proposed G-PFZF and G-PWPFZF schemes with local weights consistently outperform PFZF and PWPFZF with LSFD across moderate to high pilot lengths, confirming the advantage of adaptive per-AP combining strategies.
@@
Fig. 9 presents the 90%-likely per-user SE across different combining schemes, demonstrating significant improvements in user fairness through our decentralized optimization framework. The proposed G-PFZF scheme achieves 45% improvement in 90%-likely SE compared to conventional PFZF, while G-PWPFZF shows a 34% gain over its baseline PWPFZF. These gains confirm that our adaptive per-AP grouping strategy not only enhances the overall system capacity but fundamentally transforms user experience across the network. The gains in 90%-likely SE particularly highlight our framework’s effectiveness in improving the performance of weak users.
@@
Fig. 10 presents the 90%-likely per-user SE comparing local versus LSFD processing, revealing crucial insights about the performance-complexity trade-off in our decentralized framework. While our G-PFZF and G-PWPFZF schemes with local weights experience a modest performance reduction of approximately 10-10.7% compared to their LSFD-based counterparts, the 10% performance difference between local and LSFD processing represents a reasonable trade-off for achieving fully decentralized architecture, substantially reduced fronthaul overhead, and lower computational cost. Noticeably, our locally-weighted G-PFZF scheme outperforms the baseline PFZF with optimal LSFD by 32%, and G-PWPFZF with local weights surpasses the LSFD-enhanced PWPFZF baseline by 22.5%. These results demonstrate that the performance gains from our adaptive per-AP grouping strategy are so significant that they outweigh the benefits of optimal LSFD coordination applied to suboptimal threshold-based grouping.
@@ V Conclusion and Future Work
This paper has addressed the critical bottlenecks of sub-optimal performance, high fronthaul load, and computational cost in D-mMIMO networks by introducing a novel, decentralized decoding uplink architecture that fundamentally departs from conventional LSFD-based designs. Through the proposed local ZF framework, we have demonstrated that enabling each AP to independently determine its combining strategy via local optimization, classifying pilots as strong or weak without fixed thresholds and dynamically switching among PFZF/PWPFZF, FZF, or MR, yields significant performance improvements. The resulting generalized schemes, G-PFZF and G-PWPFZF, consistently outperform their fixed-threshold counterparts across all evaluated scenarios, achieving substantial gains in both sum and per user spectral efficiency. Moreover, the introduction of pilot-dependent combining vectors and interference-suppressing local weights applied distributively at each AP eliminates the need for centralized LSFD, thereby drastically reducing fronthaul overhead and computational cost. Remarkably, this decentralized approach incurs only a small performance penalty compared to idealized centralized coordination, while even outperforming conventional threshold-based schemes using optimal LSFD. These findings collectively establish that proposed decentralized decoding in the network, not centralized decoding, is the key to scalability and efficiency in future D-mMIMO systems without compromising on quality-of-service.
In future, we plan to investigate the robustness of the proposed decentralized architecture and generalized combining schemes under more practical and challenging conditions, such as hardware impairments.
-A Derivation of closed-form SINR expression for G-PFZF combining
In order to derive the closed-form expression, we need few expectations computation or expansions
$\begin{split}\mathbb{E}\Big\{\Big|\!\!\sum_{m}\!\!\mathbf{v}_{mi}^{H}\mathbf{g}_{mi}\!\Big|^{2}\!\Big\}\!\!=&\!\sum_{m}\!\mathbb{E}\Big\{\!\Big|\mathbf{v}_{mi}^{H}\mathbf{g}_{mi}\!\Big|^{2}\!\Big\}+\!\Big|\!\sum_{m}\!\mathbb{E}\Big\{\mathbf{v}_{mi}^{H}\mathbf{g}_{m}\Big\}\Big|^{2}-\sum_{m}\Big|\mathbb{E}\Big\{\mathbf{v}_{mi}^{H}\mathbf{g}_{mi}\Big\}\Big|^{2},\end{split}$ (A1)
$\mathbb{E}\Big\{\Big|\mathbf{v}_{mi_{t}}^{H}\tilde{\mathbf{g}}_{mk}\Big|^{2}\Big\}=$ $\beta_{mk}-\gamma_{mk}$ (A2)
$\mathbb{E}\Big\{\Big|\mathbf{v}_{mi_{t}}^{H}{\mathbf{n}}_{m}\Big|^{2}\Big\}=$ $1.$ (A3)
The desired signal term in (6) can be written as
@@
where $\mathcal{Z}_{i_{t}}$ and $\mathcal{Y}_{i_{t}}$ denote the subsets of APs where pilot $i_{t}$ is classified as strong and weak, respectively. Using (10) and (12), the $|\text{DS}_{t}|^{2}$ can be written as:
$\begin{split}|\text{DS}_{t}|^{2}&=\!\!\Big|\!\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{(A{-}L_{\mathcal{S}_{m}})\gamma_{mk}}+\!\!\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{A\gamma_{mk}}\Big|^{2}\\ &=\Big|\sum_{m=1}^{M}\!a_{mt}\sqrt{(A-\delta_{mi_{t}}L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2},\end{split}$ (A4)
@@
$\begin{split}\mathbb{E}\Bigg\{\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{g}}_{mk}+\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{g}}_{mk}\Big|^{2}\Bigg\}&=\!\mathbb{E}\Bigg\{\!\Big|\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}\Big|^{2}\!\!\Bigg\}\!+\mathbb{E}\Bigg\{\!\Big|\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{mk}\Big|^{2}\!\!\Bigg\}\\ &+2\mathbb{E}\Bigg\{\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!\sum_{\ \ \ n\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!a_{mt}a_{nt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}(\mathbf{v}_{ni_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{nk}\!\Bigg\}.\end{split}$ (A5)
Using (A1), the first term ($A_{1}$) of (A5) can be rewritten as:

### v3

@@ IV-B Numerical Simulations
Fig. 4 shows that the baseline PFZF and PWPFZF schemes exhibit a pronounced peak at $L_{\mathcal{S}_{m}}{=}6$-77, revealing their rigid threshold-based strategy that sacrifices excessive degrees of freedom for interference suppression regardless of local conditions. In striking contrast, the proposed G-PFZF and G-PWPFZF schemes show fundamentally different behavior: G-PFZF and G-PWPFZF peaks at 1-4 strong pilots. This efficient allocation preserves spatial resources for signal enhancement. Most notably, the significant non-zero count at $L_{\mathcal{S}_{m}}{=}0$ demonstrates our algorithm’s ability to adaptively default to simple MR combining when it is the optimal strategy for a given AP’s local conditions. This demonstrates a seamless, adaptive switching between ZF and MR combining.
Figure 5: The uplink sum SE comparison for various decoding schemes.
Figure 6: The uplink sum SE comparison for various decoding schemes.
Figs. 5 and 6 illustrate the sum SE versus the number of UEs for the proposed G-PFZF and G-PWPFZF combining schemes, respectively, comparing c-LSFD, the proposed d-LSFD, and h-LSFD [13]. For both G-PFZF and G-PWPFZF, the performance gap between d-LSFD and c-LSFD remains consistently within 1–2% across all UE densities. This confirms that the proposed locally computable LSFD weights achieve near-optimal performance within the class of decentralized linear decoders, while completely eliminating the need for network-wide coordination and centralized matrix inversions. In contrast, h-LSFD exhibits a substantially larger performance loss of approximately 10–12%, highlighting the benefit of the proposed principled weight design.
Fig. 6 additionally includes LP-MMSE combining for comparison. While c-LSFD achieves the highest performance for LP-MMSE, the gap between c-LSFD and d-LSFD is very significant. This behavior is consistent with the theoretical analysis in Fig. 3, which shows that the SINR guarantee for d-LSFD becomes vacuous when the $\rho_{t}$ is large, as is typical for MMSE-based local combining due to strong global interference coupling. These results further validate that the proposed d-LSFD scheme is particularly effective for ZF-type combining schemes.
Figure 7: The uplink sum SE comparison for various decoding schemes.
Fig. 7 illustrates the sum SE versus the number of pilots ($L_{p}$), comparing decoding schemes for G-PFZF and G-PWPFZF combining. For G-PFZF and G-PWPFZF combining, the performance gap between d-LSFD and c-LSFD is marginal (less than 3%), when $L_{p}=5$, as high pilot contamination makes c-LSFD better equipped to handle interference. However, as the number of pilots increases to moderate values, the gap reduces to negligible. This trend occurs because with sufficient orthogonal pilots, both approaches have similar capability to mitigate pilot contamination. Notably, even in the worst-case scenario with limited number of orthogonal pilots, the performance penalty remains modest, validating that the substantial fronthaul and computation reductions of our d-LSFD approach come with acceptable performance trade-offs across all pilot sequences. Furthermore, the proposed d-LSFD outperform the h-LSFD with a significant margin across all pilot densities for both the combining schemes. These results further show that the proposed d-LSFD retains near-centralized performance, for all pilot densities.
@@
Fig. 8 depicts the sum SE versus the number of UEs ($T$) for various combining schemes. Compared with the baseline PFZF and PWPFZF schemes, when using either c-LSFD or d-LSFD, the proposed G-PFZF and G-PWPFZF with d-LSFD achieve gains of 6%-9.5% and 5%-10%, respectively, with the gaps widening as the number of UEs increases. This performance improvement comes from the adaptive per-AP pilot classification, which replaces the fixed threshold-based UE grouping of baselines. This enables a more efficient and dynamic trade-off between utilizing degrees of freedom for interference suppression and desired signal enhancement. Consequently, the proposed G-PWPFZF scheme achieves better performance over the G-PFZF scheme by suppressing interference of the weak UEs from the strong UEs at the APs. By comparing PFZF and PWPFZF under both c-LSFD and d-LSFD with their proposed generalized counterparts, the observed performance gains persist irrespective of the decoding strategy. This confirms that the improvements primarily stem from the proposed adaptive pilot-based combining enabled by local optimization.
@@
Fig. 9 illustrates the sum SE versus the number of orthogonal pilots $L_{p}$, demonstrating the improved performance of the proposed G-PFZF and G-PWPFZF with d-LSFD schemes over their respective baseline counterparts with c-LSFD and d-LSFD. The proposed schemes achieve significant performance improvements of 6-14% for G-PFZF and 5-13% for G-PWPFZF compared to their respective baselines across different pilot lengths, with the performance gap widening as the number of pilots increases. This expanding gap highlights the advantage of the proposed adaptive pilot grouping strategy over the rigid network-wide threshold approach used in baseline schemes. The baseline schemes must rigidly limit the number of strong pilots to avoid violating the fundamental condition for zero-forcing ($L_{\mathcal{S}_{m}}{>}A$), which leads to inefficient interference management. In contrast, our proposed decentralized optimization, inherently ensures $L_{\mathcal{S}_{m}}{<}A$, while maximizing array gain and suppressing interference.
@@
Fig. 10 presents the 90%-likely per-user SE across different combining schemes, demonstrating significant improvements in per user SE through our decentralized optimization framework. The proposed G-PFZF with d-LSFD scheme achieves 40-45% improvement in 90%-likely SE compared to conventional PFZF with both decoding scheme, while G-PWPFZF with d-LSFD shows a 28-34% gain over its baseline PWPFZF across both decoding scheme. These gains confirm that our adaptive per-AP grouping strategy not only enhances the overall system capacity but fundamentally transforms user experience across the network. The gains in 90%-likely SE, particularly, highlight our optimization framework’s effectiveness in improving the performance of weak users.
@@
Fig. 11 presents the 90%-likely per-user SE. While our proposed d-LSFD experience a modest performance reduction of approximately 5.5% and 7% compared to the c-LSFD for G-PFZF and G-PWPFZF scheme respectively. The less than 7% performance difference between d-LSFD and c-LSFD represents a reasonable trade-off for achieving fully decentralized architecture, substantially reduced fronthaul overhead, and lower computational cost. Noticeably, our proposed d-LSFD outperforms the h-LSFD by 100% and 65% for G-PFZF and G-PWPFZF combining respectively, thus outperforming the h-LSFD with significant margin. Overall, the results confirm that near-optimal decentralized decoding and adaptive combining address two orthogonal limitations of existing D-mMIMO systems and jointly enable scalable, high-performance uplink processing.
@@ V Conclusion and Future Work
This paper establishes a principled framework for uplink processing in D-mMIMO systems by addressing both local combining and decentralized decoding. We propose two adaptive pilot-aware local combining schemes, G-PFZF and G-PWPFZF, where pilots are partitioned using local optimization. Furthermore, we establish a universal SINR lower bound that rigorously characterizes the performance loss incurred by decentralization of LSFD. A key insight from our theoretical analysis is that the effectiveness of decentralization hinges on the structure of the local combiner. While the proposed d-LSFD is not universally optimal for all combiners, we demonstrate that for the class of pilot-aware ZF-based scheme, it attains near-centralized performance. Numerical results confirmed that, for G-PFZF and G-PWPFZF combining, d-LSFD attains near-optimal performance. In contrast, for LP-MMSE combining, d-LSFD exhibits a larger performance gap due to the presence of strong global interference coupling. Overall, the proposed framework provides a scalable and theoretically grounded alternative to c-LSFD. In future, we plan to investigate the robustness of generalized combining schemes and the decentralized decoding under hardware impairments.
-A Proof of Lemma 1
On the domain where $F(x,q)$ is well defined, the denominator is strictly positive. Hence, minimizing $F(x,q)$ over $|x|^{2}\leq q$ is equivalent to maximizing
$G(x)\triangleq\frac{q-|x|^{2}}{|1+x|^{2}}.$
Write $x=r(\cos\theta+i\sin\theta)$ with $0\leq r\leq\sqrt{q}$. Then
$G(r)=\frac{q-r^{2}}{1+2r\cos\theta+r^{2}}.$
For fixed $r$, the $G(r)$ is minimized when $\cos\theta=-1$, yielding
$G(r)=\frac{q-r^{2}}{(1-r)^{2}}.$
For $0{\leq}q{<}1$, the stationary point r=qr{=}q lies in the admissible interval $0,\sqrt{q}]$ and gives $G(q){=}\frac{q}{1-q}$. Substituting into F=1/(1+G)F{=}1/(1+G) gives
$\min_{|x|^{2}\leq q}F(x,q)=1-q.$
For $q\geq 1$, $G(r)$ can be made arbitrarily large, and therefore
$\inf_{|x|^{2}\leq q}F(x,q)=0.$
-B Proof of Theorem 2
$\text{We know,}\ \ \mathrm{SINR}_{t}(\mathbf{a}_{t}^{\mathrm{o}})=\frac{|(\mathbf{a}_{t}^{\mathrm{o}})^{H}\mathbf{d}_{t}|^{2}}{(\mathbf{a}_{t}^{\mathrm{o}})^{H}\mathbf{J}_{t}\mathbf{a}_{t}^{\mathrm{o}}}=(\mathbf{a}_{t}^{\mathrm{o}})^{H}\mathbf{J}_{t}\mathbf{a}_{t}^{\mathrm{o}}>0.$
Define the error vector $\bm{\delta}_{t}\triangleq\mathbf{a}_{t}^{\mathrm{d}}-\mathbf{a}_{t}^{\mathrm{o}}$, so that $\mathbf{a}_{t}^{\mathrm{d}}=\mathbf{a}_{t}^{\mathrm{o}}+\bm{\delta}_{t}$ and also $\|\bm{\delta}_{t}\|_{2}=\rho_{t}\|\mathbf{a}_{t}^{\mathrm{o}}\|_{2}$ Using $\mathbf{d}_{t}=\mathbf{J}_{t}\mathbf{a}_{t}^{\mathrm{o}}$, define the normalized quantities $x\triangleq\frac{\bm{\delta}_{t}^{H}\mathbf{d}_{t}}{(\mathbf{a}_{t}^{\mathrm{o}})^{H}\mathbf{J}_{t}\mathbf{a}_{t}^{\mathrm{o}}}$, and $q\triangleq\frac{\bm{\delta}_{t}^{H}\mathbf{J}_{t}\bm{\delta}_{t}}{(\mathbf{a}_{t}^{\mathrm{o}})^{H}\mathbf{J}_{t}\mathbf{a}_{t}^{\mathrm{o}}}$. A direct expansion yields
$\frac{\mathrm{SINR}_{t}(\mathbf{a}_{t}^{\mathrm{d}})}{\mathrm{SINR}_{t}(\mathbf{a}_{t}^{\mathrm{o}})}=\frac{|1+x|^{2}}{1+2\Re(x)+q}.$ (T1)
By the Rayleigh quotient bounds,
$q\leq\frac{\lambda_{\max}(\mathbf{J}_{t})\|\bm{\delta}_{t}\|_{2}^{2}}{\lambda_{\min}(\mathbf{J}_{t})\|\mathbf{a}_{t}^{\mathrm{o}}\|_{2}^{2}}=\kappa_{t}\rho_{t}^{2},$
and by the Cauchy-Schwarz inequality in the $\mathbf{J}_{t}$-inner product, $|x|^{2}\leq q$. Using Lemma 1 for fixed $q$, we yields
$\min_{|x|^{2}\leq q}\frac{|1+x|^{2}}{1+2\Re(x)+q}=\max\{0,\;1-q\}.$
Substituting $q\leq\kappa_{t}\rho_{t}^{2}$ completes the proof.
-C Proof of Proposition 1
Consider an off-diagonal entry of $\mathbf{J}_{t}$ for $m\neq n$:
$[\mathbf{J}_{t}]_{mn}$ $=\sum_{k=1,k\neq t}^{T}p_{k}^{u}\mathbb{E}\!\Big\{\mathbf{v}_{mt}^{H}\hat{\mathbf{g}_{mk}}\Big\}\mathbb{E}\!\Big\{\mathbf{v}_{nt}^{H}\hat{\mathbf{g}}_{nk}\Big\}.$
The local MR and ZF combining vectors are linearly dependent only on the channel estimates of UE $t$ and the UEs sharing the same pilot as $t$, since pilot-sharing UEs have linearly dependent channel estimates as in (1). Hence, for any $k\neq t$ that does not share pilot with $t$, $\mathbb{E}\!\left\{\mathbf{v}_{mt}^{H}\hat{\mathbf{g}}_{mk}\right\}=0$, which implies $\mathbf{J}_{t}]_{mn}=0$ for all $m\neq n$. Therefore, $\mathbf{J}_{t}$ becomes a diagonal matrix.
-D Closed-form SINR and d-LSFD expressions for G-PFZF combining
Pilot-sharing UEs use $\mathbf{v}_{mi_{t}}^{\mathrm{LZF}}$ for strong pilots and $\mathbf{v}_{mi_{t}}^{\mathrm{MR}}$ for weak pilots. The desired signal term in (7) can be written as
@@
where $\mathcal{Z}_{i_{t}}\subseteq\mathcal{M}_{t}$ and $\mathcal{Y}_{i_{t}}\subseteq\mathcal{M}_{t}$ denote the subsets of APs where pilot $i_{t}$ is classified as strong and weak, respectively. Using (12) and (14), the $|\text{DS}_{t}|^{2}$ can be written as:
$\begin{split}|\text{DS}_{t}|^{2}&=\!\!\Big|\!\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{(A{-}L_{\mathcal{S}_{m}})\gamma_{mt}}+\!\!\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{A\gamma_{mt}}\Big|^{2}\\ &=\Big|\sum_{m=1}^{M}\!a_{mt}\sqrt{(A-\delta_{mi_{t}}L_{\mathcal{S}_{m}})\gamma_{mt}}\Big|^{2},\end{split}$ (A1)
@@
$\begin{split}&\mathbb{E}\Big\{\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{g}}_{mk}+\sum_{m\in\mathcal{Y}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{g}}_{mk}\Big|^{2}\Big\}\\ &=\!\mathbb{E}\Big\{\!\Big|\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}\Big|^{2}\!\Big\}\!+\mathbb{E}\Big\{\!\Big|\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{mk}\Big|^{2}\!\Big\}\\ &+2\mathbb{E}\Big\{\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!\sum_{\ \ \ n\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!a_{mt}a_{nt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}(\mathbf{v}_{ni_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{nk}\!\Big\}.\end{split}$ (A2)
The first term ($A_{1}$) of (A2) can be rewritten as: